OpenStudy (anonymous):

I'm not sure how this one is done because of the fraction but the problem just says "add (-9x+12)+(2x^2+1/2x-14)" and when I try to solve I get the answer 5&1/2x^2-14 but it seems like that is wrong for some reason

OpenStudy (mertsj):

\[-9x+12+2x^2+\frac{1}{2}x-14=2x^2+(-9x+\frac{1}{2}x)+(12-14)\]

OpenStudy (mertsj):

Did you notice that I grouped the like terms?

OpenStudy (anonymous):

yes, I tried that too but failed :(

OpenStudy (mertsj):

\[2x^2-\frac{17}{2}x-2\]

OpenStudy (anonymous):

how did you get 17 for the numerator?

OpenStudy (mertsj):

-9+1/2=-8 1/2 = -17/2

OpenStudy (anonymous):

oh duh I see now lol thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

but I think I am supposed to simplify so would it be the -8 1/2 then right?

OpenStudy (anonymous):

yes u have to simplify

OpenStudy (mertsj):

It's hard to say what your teacher requires. In lower grades, the mixed number is considered simple but in algebra and above, the improper fraction is usually considered fine.

OpenStudy (anonymous):

sorry lol I know I take patience to try to teach math

OpenStudy (mertsj):

Maybe you could put it both ways

OpenStudy (anonymous):

well technically both are correct, I will try that, thanks again!

OpenStudy (mertsj):

yw